# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

### v1

@@ 2.2 Ideal Utility and Benchmarks
If each agent $i$ has value distribution $\mathcal{F}_{i}=\mathrm{Bernoulli}(\alpha_{i})$, it is impossible to guarantee each agent $i$ a $\lambda$ fraction of their ideal utility in expectation for $\lambda>1-\prod_{j\in[n]}(1-\alpha_{j})$, even if the mechanism knows $V_{i}[t]$ for all $i,t$ before round $1$. Note that $\inf_{n,\alpha_{1},\ldots,\alpha_{n}}1-\prod_{j\in[n]}(1-\alpha_{j})=1-1/e$.
Proof.
$\frac{1}{T}\sum_{t=1}^{T}\sum_{i\in[n]}\mathbb{E}[U_{i}[t]]\leq 1-\prod_{j\in[n]}(1-\alpha_{j}).$ (1)
With these Bernoulli value distributions, each agent’s ideal utility is exactly $\alpha_{i}$. If each agent is obtaining a $\lambda$ fraction of their ideal utility in expectation, then
$\frac{1}{T}\sum_{t=1}^{T}\sum_{i\in[n]}\mathbb{E}[U_{i}[t]]\geq\lambda\sum_{i\in[n]}\alpha_{i}=\lambda,$
so $\lambda\leq 1-\prod_{j\in[n]}(1-\alpha_{j})$ by (1). ∎
3 Warm Up: Repeated Allocation with Two Agents
To motivate the design of our mechanism and show how we can achieve good robust equilibria as outlined in Section 2.2, we first give a simplified version of our mechanism in the special case where there are only two agents. Our mechanism lets each agent decide to bid for the resource or not in each round, up to some limited number of bids. It makes sense that when only one agent bids (and has not exceeded her bid budget), she is guaranteed the item. What is unclear is who gets allocated when both agents bid. We show that carefully randomizing this allocation leads to the optimal guarantee of Lemma 2.1, both in terms of the equilibrium and the robust setting, i.e., $\lambda_{\mathtt{NASH}}=\lambda_{\mathtt{ROB}}=1-(1-\alpha_{1})(1-\alpha_{2})$. To extend this result to $n$ agents, we need to set exponentially many parameters (the allocations when any subset $S$ agents bids), which we do by strengthening Border’s theorem in Section 5.
Our mechanism proceeds as follows: Each agent $i$ starts with $\alpha_{i}T$ bid tokens. At each round, each agent can bid 00 or $1$. If both agents bid 00, no one receives the item. If only one agent bids $1$, then she receives the item in that round. If both agents bid $1$, then we allocate the item to agent $1$ with probability $p_{1}^{\{1,2\}}=p$ and to agent $2$ with probability $p_{2}^{\{1,2\}}=1-p$. Regardless of who wins the item, each agent $i$ pays their bid. If an agent’s budget becomes 00, they can not bid.
Proposition 3.1 (Informal).
Proof sketch.
First, observe that the expected spending of each agent $i$ under this strategy profile is indeed at most $B_{i}[1]=\alpha_{i}T$, since each agent will be bidding i.i.d. $\mathrm{Bernoulli}(\alpha_{i})$. Next, note that there is nothing that agent $1$ can do to affect player $2$’s behavior. She must solve her own stochastic control problem where her policy does not affect agent $2$’s behavior. At any time $t$, if agent $1$ bids, her probability of winning is $p_{1}=\alpha+(1-\alpha)p$, where the $\alpha$ term is the probability that agent $2$ does not bid, and the $(1-\alpha)p$ term is the probability that agent $2$ bid, and agent $1$ wins give both agents bid. Thus, agent $1$’s control problem reduces to selecting $\alpha_{1}T$ time periods in which to bid, and she wins a $p_{1}$ fraction of them in expectation. Agent $1$ should bid on her highest-valued $\alpha_{i}T$ time periods, which is precisely what bidding whenever her value $V_{i}[t]$ is in the top $\alpha_{i}$-quantile of her distribution does (in expectation). ∎
where $p_{1}=\alpha+(1-\alpha)p$ is the probability agent $1$ wins a round conditioned on her bidding. Symmetrically, player $2$ gets expected utility $\frac{1}{T}\sum_{t}\mathbb{E}[U_{2}[t]]=p_{2}v_{2}^{\star}$ where $p_{2}=(1-\alpha)+\alpha(1-p)=1-\alpha p$. Setting $p=1-\alpha$ guarantees both agents $1-\alpha(1-\alpha)$ fraction of their ideal utility. This achieves $\lambda_{\mathtt{NASH}}=1-\alpha(1-\alpha)$ and by Lemma 2.1, this is the best factor any allocation rule can get.
The choice of $p=1-\alpha$ might seem unusual, especially when $\alpha$ is close to 00 or $1$ where one agent wins most times. The intuition behind this is as follows: if agent $1$ has small fair share and agent $2$ has a high one, then using a high $p$ and favoring agent $1$ has a small effect on agent $2$, most of whose bids are uncontested. Therefore, if we want equal outcomes, i.e., $p_{1}=p_{2}$, we have to favor the agent with the small fair share, breaking most ties in her favor.
Proposition 3.2 (Informal).
This means we achieve $\lambda_{\mathtt{ROB}}=1-(1-\alpha_{1})(1-\alpha_{2})$ which also matches the upper bound of any allocation of Lemma 2.1.
In coming sections, we generalize this mechanism for $n\geq 2$ players, where the main difficulty is resolving concurrent bids for any subset $S\subseteq[n]$ with $|S|\geq 2$. Since there are exponentially many such subsets, it is much harder to find a good or simple allocation rule that has the same properties as our choice of $p$ above. We will show that we can get the same optimal factor under equilibrium performance, $\lambda_{\mathtt{NASH}}=1-\prod_{i}(1-\alpha_{i})$, using Border’s theorem. For the robust guarantee, we show that a simple application of Border’s theorem is not enough. Instead, we have to consider more restrictions on how we handle concurrent bids, which leads to the strengthening of Border’s theorem in Lemma 6.2. Unfortunately, the robust guarantee weakens for $n>2$, and we get $\lambda_{\mathtt{ROB}}\approx 1/2$, which we show is tight under any rule for resolving concurrent bids.
@@ 4 Mechanism and Proposed Strategy
Input: Fair shares $\alpha_{i}$, number of rounds $T$, allocation probabilities $(p_{i}^{S})_{i\in S},\forall S\subseteq[n]$, budget slack $\delta_{i}^{T}$.
Endow each agent with a budget $B_{i}[1]=\alpha_{i}(1+\delta_{i}^{T})T$ of bid tokens;
@@
      Agents submit bids $b_{i}^{t}\in\{0,1\}$;
      Budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $B_{i}[t]\leq 0$;
      Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents;
      A winner $i^{t}$ is randomly selected from $S[t]$ according to a probability distribution $(p_{i}^{S[t]})_{i\in S[t]}$;
      Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$;
@@
MECHANISM 1 Budgeted Robust Border (BRB) Mechanism
We now extend our mechanism from Section 3 and present it for the general case of $n\geq 2$. Similar to Section 3, every agent $i$ bids for the item or not in every round and can make (approximately) at most $\alpha_{i}T$ bids in total. In every round where agents bid for the resource, our mechanism randomly allocates the resources to one of those agents. Specifically, if agents $S\subseteq[n]$ bid for the resource, we allocate the resource to agent $i\in S$ with probability $p_{i}^{S}$. We call the $p_{i}^{S}$’s allocation probabilities and describe their values later (Theorem 4.2). Our formal mechanism can be found in Mechanism 1, where we also provide a little extra budget for each player, to ensure they do not run out with high probability when following our proposed equilibrium strategy, which we introduce next.
The strategy class we propose is a simple and natural strategy wherein each agent bids whenever their value is above a certain threshold. Adopting terminology from [fikioris2023online], we refer to this as a $\beta$-aggressive strategy.
@@
Intuitively, for an agent with fair share $\alpha$, following a $\alpha$-aggressive strategy is akin to truthfully reporting which are her $\alpha T$ favorite rounds. Building on this, we propose that each agent $i$ follows an $\alpha_{i}$-aggressive strategy. Note that with no competition, this would realize the agent’s fair share. Our first result is that everyone playing an $\alpha_{i}$-aggressive strategy is an approximate Nash equilibrium.
@@
By setting $\delta_{i}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{i}T}}}$, no matter the choice of allocation probabilities $p_{i}^{S}$, each agent playing an $\alpha_{i}$-aggressive strategy is an $O(\sqrt{\nicefrac{{\log T}}{{T}}})$-approximate Nash equilibrium.
We defer the proof of this to Section 5. Note that the utility guarantees at this equilibrium depend on the choice of allocation probabilities $p_{i}^{S}$. Since the agents’ bids are i.i.d. in every round, we can calculate the fixed probability $p_{i}$ that agent $i$ will win the item conditioned on bidding (assuming everyone has budget remaining). This $p_{i}$ is called agent $i$’s interim allocation probability.
@@
Our goal is to maximize the fraction of ideal utility every agent is guaranteed in the above equilibrium. By definition of $p_{i}$, every agent in the equilibrium is guaranteed a $p_{i}$ fraction of their ideal utility. The following theorem shows that we can set $p_{i}^{S}$ to achieve the upper bound of Lemma 2.1; we present the proof (which is based on Border’s theorem) in Section 5.
@@
There exist allocation probabilities $p_{i}^{S}$ such that the induced interim allocation probabilities $p_{i}$ are the same for all agents $i$. Specifically, we can make
$\forall i:\quad p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j}).$
Having achieved optimal equilibrium performance, we turn our attention to robustness. Unlike the equilibrium analysis, where we can use any allocation probabilities $p_{i}^{S}$ that induce the desired interim allocation probabilities $p_{i}$, under arbitrary competition, we must be more careful with our selection of $p_{i}^{S}$. In Section 6.2, we show that not carefully picking the $p_{i}^{S}$’s can lead to very poor robustness. Our main technical contribution is in Section 6.1, where we show that under careful upper bounds on the $p_{i}^{S}$’s we can guarantee strong robust performance without changing the $p_{i}$’s and thus the equilibrium performance. The following theorem summarizes our overall guarantees, subsuming the results we presented in this section, and the ones we prove in the next.
@@
Consider the BRB Mechanism (Mechanism 1) with $\delta_{i}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{i}T}}}$. With a careful choice of allocation probabilities $p_{i}^{S}$ satisfying Theorem 4.2, we have the following:
@@
    Each player $i$ playing an $\alpha_{i}$-aggressive strategy is an $O(\sqrt{\nicefrac{{\log T}}{{T}}})$ Bayes-Nash equilibrium where, with probability $1-O(1/T^{2})$, player $i$ realizes utility
@@
    Regardless of behavior of others, playing an $\alpha_{i}$-aggressive strategy, with probability $1-O(1/T^{2})$, gives player $i$ utility
@@
Note that this implies that each player gets a $\lambda_{\mathtt{NASH}}=(1-\prod_{j=1}^{n}(1-\alpha_{j}))-O(\sqrt{\nicefrac{{\log T}}{{T}}})$ fraction of their ideal utility in expectation at the approximate equilibrium, and the $\alpha_{i}$-aggressive strategy is $\lambda_{\mathtt{ROB}}=(\frac{1}{2}-O(\sqrt{\nicefrac{{\log T}}{{T}}}))$-robust. The robustness factor of $1/2$ matches the best current robustness factors [gorokh2019remarkable, fikioris2023online]  [^2] Through personal communication, we have learned of ongoing work proposing a different mechanism that achieves a higher robustness factor. However, this requires a much larger strategy space, and the robust strategy is more complicated.. We prove the equilibrium claim in Section 5 and the robustness claim in Section 6.
By our choice of $\delta_{i}^{T}$ in Theorem 4.3, the probability of any agent running out of budget is $O(1/T^{2})$ (similar to the probability of our bound not holding). Due to this, we could modify BRB’s budget constraint to allow agent $i$ to bid at most $\alpha_{i}(1+\delta_{i}^{t})t$ times by round $t\in[T]$ to get any time guarantees. Specifically, our utility bounds would hold for every round $t\in[T]$ with probability $1-O(1/\sqrt{t})$. Enforcing the budget constraint at each time also allows us to obtain an exact Nash equilibrium in the infinite time horizon. See Appendix B for details.
@@ 5 Equilibrium and Good Allocation Probabilities
5.1 Equilibrium
In this section, we prove our claim of Section 4 that each agent $i$ following the $\alpha_{i}$-aggressive strategy in the BRB Mechanism forms an approximate Nash Equilibrium. The proof sketch is very similar to the one for two agents in Section 3. As in Section 3, we consider a simpler game where budgets are enforced in expectation. While bounds in expectation are not enforceable, we focus on this case for simplicity of presentation, and only require that $\sum_{t=1}^{T}\mathbb{E}[b_{i}^{t}]\leq\alpha_{i}T$. Roughly speaking, we prove that in this game, the best response of agent $i$ when agents $j\neq i$ are following an $\alpha_{j}$-aggressive strategy is to follow an $\alpha_{i}$-aggressive strategy also. This, in turn, proves that these strategies form an equilibrium.
@@ 5.1 Equilibrium
Fix an agent $i$ and assume all other agents $j\neq i$ are bidding i.i.d. $\mathrm{Bernoulli}(\alpha_{j})$. Suppose agent $i$ is trying to maximize her expected utility subject to the constraint that she does spend more than $\alpha_{i}T$ tokens in expectation; that is, she is choosing $(b_{i}^{t})_{t}$ to solve
@@
Her optimal strategy is to choose $(b_{i}^{t})_{t}$ that corresponds to bidding whenever her value is in the top $\alpha_{i}$-quantile of her value distribution, which yields agent $i$ expected utility
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]=p_{i}v_{i}^{\star}.$
The idea behind why the $\alpha_{i}$-aggressive strategy is the best response is as follows. At each round, the probability that agent $i$ wins conditioned on bidding is exactly their interim allocation probability $p_{i}$. The best solution for the agent is to bid in the $\alpha_{i}T$ rounds where the agent has the highest value. Therefore, agent $i$ should follow an $\alpha_{i}$-aggressive strategy, bidding whenever her value is in the top $\alpha_{i}$-quantile of her value distribution. We include the full proof of the lemma in Appendix D, where we also translate the result from this simplified game with budget constraints met in expectation to the actual game where budgets are a bit larger and strictly enforced.

### v2

@@ 2.1 Ideal Utility and Benchmarks
If each agent $i$ has value distribution $\mathcal{F}_{i}=\mathrm{Bernoulli}(\alpha_{i})$, it is impossible to guarantee every agent $i$ a $\lambda$ fraction of their ideal utility in expectation for $\lambda>1-\prod_{j\in[n]}(1-\alpha_{j})$, even if the mechanism knows $V_{i}[t]$ for all $i,t$ before round $1$. Note that $\inf_{n,\alpha_{1},\ldots,\alpha_{n}}1-\prod_{j\in[n]}(1-\alpha_{j})=1-1/e$.
We prove the lemma in subsection˜E.1.
@@ 3 Mechanism and Proposed Strategy
We now give our mechanism. Every agent $i$ bids for the item or not in every round and can make (approximately) at most $\alpha_{i}T$ bids in total. In every round where agents bid for the resource, our mechanism randomly allocates the resources to one of those agents. For the simpler randomized allocation rule in the case where there are only $2$ agents, see section˜A [^1] When $n=2$, in section A, we manage to get optimal performance $\lambda_{\mathtt{NASH}}=\lambda_{\mathtt{ROB}}=1-(1-\alpha_{1})(1-\alpha_{2})$, as in Lemma 2.2.. In general, if agents $S\subseteq[n]$ bid for the resource, we allocate the resource to agent $i\in S$ with probability $p_{i}^{S}$. We call the $p_{i}^{S}$’s allocation probabilities and describe their values later (Theorem˜3.5). Our formal mechanism can be found in Mechanism˜1, where we also provide a little extra budget for each agent, to ensure they do not run out with high probability when following our proposed equilibrium strategy, which we introduce next.
Mechanism 1 Budgeted Robust Border (BRB) Mechanism
0: Fair shares $\alpha_{i}$, time horizon $T$, allocation probabilities $(p_{i}^{S})_{i\in S}\;\forall S\subseteq[n]$, budget slack $\delta_{i}^{T}$.
 Endow each agent $i$ with a budget $B_{i}[1]=\alpha_{i}(1+\delta_{i}^{T})T$ of bid tokens.
@@
  Agents submit bids $b_{i}^{t}\in\{0,1\}$ and budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $B_{i}[t]\leq 0$.
  Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents.
  A winner $i^{t}$ is randomly selected from $S[t]$ according to the probability distribution $(p_{i}^{S[t]})_{i\in S[t]}$.
  Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$.
@@
The strategy class we propose is a simple and natural strategy wherein each agent bids whenever their value is above a certain threshold. As in [fikioris2023online], we refer to this as a $\beta$-aggressive strategy.
@@
Intuitively, for an agent with fair share $\alpha$, following a $\alpha$-aggressive strategy is akin to truthfully reporting her $\alpha T$ favorite rounds. Building on this, we propose that each agent $i$ follows an $\alpha_{i}$-aggressive strategy. Note that with no competition, this would realize the agent’s ideal utility. Our first result is that everyone playing an $\alpha_{i}$-aggressive strategy is an approximate Nash equilibrium for any allocation probabilities.
Definition 3.2 ($\varepsilon$-approximate Nash equilibrium).
For a profile of strategies $\pi=(\pi_{1},\dots,\pi_{n})$, let $\mathcal{U}_{i}(\pi)$ denote the per-rounded expected utility $\frac{1}{T}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[U_{i}[t]]$ of agent $i$ when players play according to $\pi$. We say that a profile of strategies $\pi=(\pi_{1},\dots,\pi_{n})$ forms an $\varepsilon$-approximate Nash equilibrium if for each agent $i$, for every profile of strategies $\pi^{\prime}=(\pi_{1}^{\prime},\dots,\pi_{n}^{\prime})$ where $\pi_{j}=\pi_{j}^{\prime}$ for each $j\neq i$, we have $\mathcal{U}_{i}(\pi^{\prime})\leq\mathcal{U}_{i}(\pi)+\varepsilon$.
@@
By setting $\delta_{i}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{i}T}}}$, for any fixed allocation probabilities $p_{i}^{S}$, each agent playing an $\alpha_{i}$-aggressive strategy is an $O(\sqrt{\nicefrac{{\log T}}{{T}}})$-approximate Nash equilibrium.
We defer the proof of this to section˜4. Note that the utility guarantees at this equilibrium depend on the choice of allocation probabilities $p_{i}^{S}$. Since the agents’ bids are i.i.d. in every round, we can calculate the fixed probability $p_{i}$ that agent $i$ will win the item conditioned on bidding (assuming everyone has budget remaining). This $p_{i}$ is called agent $i$’s interim allocation probability.
@@
Our goal is to maximize the fraction of ideal utility every agent is guaranteed in the above equilibrium. By definition of $p_{i}$, every agent in the equilibrium is guaranteed a $p_{i}$ fraction of their ideal utility. The following theorem shows that we can set $p_{i}^{S}$ to achieve the upper bound of Lemma˜2.2; we explain the theorem in more detail in terms of Border’s theorem in section˜4.
@@
There exist allocation probabilities $p_{i}^{S}$ such that the induced interim allocation probabilities $p_{i}$ are the same for all agents $i$. Specifically, we can make $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ for all $i$, matching Lemma˜2.2.
Having achieved optimal equilibrium performance, we turn our attention to robustness. Unlike the equilibrium analysis, where we can use any allocation probabilities $p_{i}^{S}$ that induce the desired interim allocation probabilities $p_{i}$, under arbitrary competition, we must be more careful with our selection of $p_{i}^{S}$. In subsection˜5.2, we show that not carefully picking the $p_{i}^{S}$’s can lead to very poor robustness. Our main technical contribution is in subsection˜5.1, where we show that under careful upper bounds on the $p_{i}^{S}$’s we can guarantee strong robust performance without changing the $p_{i}$’s and thus the equilibrium performance. The following theorem summarizes our overall guarantees, subsuming Theorem˜3.5.
@@
Consider the BRB Mechanism (Mechanism˜1) with $\delta_{i}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{i}T}}}$. With a careful choice of allocation probabilities $p_{i}^{S}$ (that satisfy Theorem˜3.5), we have the following:
@@
    Each agent $i$ playing an $\alpha_{i}$-aggressive strategy is an $O(\sqrt{\nicefrac{{\log T}}{{T}}})$-approximate Bayes-Nash equilibrium where, with probability $1-O(1/T^{2})$, agent $i$ realizes utility
@@
    Regardless of behavior of others, playing an $\alpha_{i}$-aggressive strategy, with probability $1-O(1/T^{2})$, gives agent $i$ utility
@@
Note that this implies that with high probability, each agent gets a $\lambda_{\mathtt{NASH}}=1-\prod_{j=1}^{n}(1-\alpha_{j})-O(\sqrt{\nicefrac{{\log T}}{{T}}})$ fraction of their ideal utility at the approximate equilibrium, and a $\lambda_{\mathtt{ROB}}=(\frac{1}{2}-O(\sqrt{\nicefrac{{\log T}}{{T}}}))$ fraction robustly. We present in more detail the equilibrium claim in section˜4 and the robustness claim in section˜5.
In sections˜4 and 5 we only show the existence of allocation probabilities $p_{i}^{S}$ and not how to efficiently compute them. However, using techniques from [bhalgat2013optimal], we give an algorithm that, given a bidding set $S$, can efficiently sample from a distribution $(p_{i}^{S})_{i\in S}$, where $((p_{i}^{S})_{i\in S})_{S\subseteq[n]}$ are allocation probabilities that will give us the equilibrium and robustness guarantees of Theorem˜3.6. We detail this algorithm and how it can be used with BRB in section˜6.
By our choice of $\delta_{i}^{T}$ in Theorem˜3.6, the probability of any agent running out of budget is $O(1/T^{2})$ (similar to the probability of our bound not holding). Due to this, we could modify BRB’s budget constraint to allow agent $i$ to bid at most $\alpha_{i}(1+\delta_{i}^{t})t$ times by round $t\in[T]$ to get any time guarantees. Specifically, our utility bounds would hold for every round $t\in[T]$ with probability $1-O(1/\sqrt{t})$. Enforcing the budget constraint at each time also allows us to obtain an exact Nash equilibrium in the infinite time horizon. See section˜C for details.
@@ 4 Equilibrium and Good Allocation Probabilities
In this section, we prove our claim of section˜3 that each agent $i$ following the $\alpha_{i}$-aggressive strategy in the BRB Mechanism forms an approximate Nash Equilibrium. For simplicity, we consider a simpler game where budgets are enforced in expectation. While bounds in expectation are not enforceable, we focus on this case for simplicity of presentation, and only require that $\sum_{t=1}^{T}\mathbb{E}[b_{i}^{t}]\leq\alpha_{i}T$. Roughly speaking, we prove that in this game, the best response of agent $i$ when agents $j\neq i$ are following an $\alpha_{j}$-aggressive strategy is to follow an $\alpha_{i}$-aggressive strategy also. This proves that these strategies form an equilibrium.
@@
Fix an agent $i$ and assume all other agents $j\neq i$ are bidding i.i.d. $\mathrm{Bernoulli}(\alpha_{j})$. Suppose agent $i$ is trying to maximize her expected utility subject to the constraint that she does not spend more than $\alpha_{i}T$ tokens in expectation; that is, she is choosing $(b_{i}^{t})_{t}$ to solve
@@
Her optimal strategy is to choose $(b_{i}^{t})_{t}$ that corresponds to bidding whenever her value is in the top $\alpha_{i}$-quantile of her value distribution, which yields agent $i$ expected utility $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]=p_{i}v_{i}^{\star}.$
To see why $\alpha_{i}$-aggressiveness is the best response, note that in each round, the probability that agent $i$ wins conditioned on bidding is exactly their interim allocation probability $p_{i}$. The best solution for the agent is to bid in the $\alpha_{i}T$ rounds where the agent has the highest value. Therefore, agent $i$ should follow an $\alpha_{i}$-aggressive strategy, bidding whenever her value is in the top $\alpha_{i}$-quantile of her value distribution. The full proof is in subsection˜E.3, where we also translate the result from this simplified game with expected budget constraints to the actual game where budgets are strictly enforced (but larger).